OpenStudy (anonymous):

Is this correct ? (4 + 2v6 ) ( 4+2v6) = 40 v stands for * sqrt

OpenStudy (anonymous):

No, its not. Remember to FOIL (4 + 2v6 )(4 + 2v6) =16 + 8v6 + 8v6 + 24 =40 + 16v6 v stands for * sqrt